OpenStudy (anonymous):

Which numbers should be multiplied to obtain 1652 - 1232? A. 42 and 42 B.42 and 288 <--- My Answer C.1,764 and 1,764 D.4,202 and 4,248

OpenStudy (anonymous):

I was unsure at first if this was correct to was just wanting to make sure

Directrix (directrix):

42 * 288 = 12096 so that is not it. None of these number pairs given multiply to 420 which is the difference of these two: 1652 - 1232

Directrix (directrix):

@caominhim I think you may want to reconsider your post.

ganeshie8 (ganeshie8):

\[\large 165^2 - 123^2\]

Directrix (directrix):

I think there is more to it. I'm thinking that a factoring pattern is involved.

OpenStudy (anonymous):

oh yea im really sorry guys Ganeshie8 was rigt that's what its suppose to look like sorry that was my bad

Directrix (directrix):

Over here at http://openstudy.com/updates/53630dbfe4b0fac5fcfecc40 , @WeakMathSkills gave a direct answer of 42 and 288 but gave no explanation.

Directrix (directrix):

Oh, you didn't show the exponents. I see.

Directrix (directrix):

Neither did @deicyg on the other rendering of this problem.

Directrix (directrix):

165^2 - 123^2 = (165 + 123)*(165 - 123) --> difference of two squares pattern

Directrix (directrix):

@Anime101 Add (165 + 123) and post what you get. Do the same for (165 - 123), okay?

OpenStudy (anonymous):

ok so (165+123) I got 288 and for the other one I got 42

Directrix (directrix):

Okay, so that is resolved. It is good to take the extra time to write the exponents as exponents. Use the carat key. 165^2 - 123^2

OpenStudy (anonymous):

oh ok thank you so much that's actually all the questions (: